# Improving IR-based Bug Localization with Semantics-Driven Query Reduction

### v1

@@ 6.1 IR based Bug Localization
Bug localization has been a key area of research for decades, driven by the significant impacts and challenges of software bugs [9]. There are two primary categories of bug localization: spectra-based and Information Retrieval (IR)-based [13]. Spectra-based methods are known for their complexity and limited scalability [78, 79]. In our work, we primarily focus on Information Retrieval for bug localization.
@@
Traditional IR-based bug localization methods [19, 18, 20, 80, 11] typically rely on the vector space model (VSM)[81], which analyzes the token overlap between bug reports and source code to identify buggy documents. Several studies have extended VSM by integrating additional contexts, such as bug report history [82], code change history [83], or version history [26], into the IR process. For instance, Zhou et al. introduced BugLocator [18] that employs a combined score of a modified VSM score and previous bug fix history to localize bugs. The traditional VSM-based scoring [81] method tends to show bias towards longer documents. To address this, they modified the VSM-based scoring and introduced rVSM scoring, which enhances the computation of textual relevance between bug reports and code elements. BLUiR [19] captures four types of structural components from the source code (i.e., class names, method names, variable names, and comments) and two components from bug reports (i.e., bug title and bug description). It then forms pairwise combinations of these components to perform eight separate searches using Indri [67], leveraging a modified TF-IDF approach. The final localization score is computed by summing the scores obtained from these individual searches. AmaLgam [20] integrates BLUiR and BugLocator techniques, along with version history inspired by Google by analyzing historical data from version control systems. These components undergo three independent systems to generate rankings before producing the final ranked lists. AmaLgam+ [80] takes a step further by incorporating stack trace and bug reporter history, alongside AmaLgam’s contexts. It ranks files using five components and then returns a final ranked list of source documents for bug localization. Recently, an IR-based technique, PathIdea [84], utilizes bug report logs (e.g., log snippets, stack traces) for bug localization. The authors employ a static analysis tool to construct a file-level call graph and reconstruct system execution paths from the logs. To determine the suspiciousness score for each file, they combine the VSM score, a log score that emphasizes files mentioned in the logs, and a path score that highlights files in the execution path. Some IR-based bug localization methods employ more complex mechanisms, such as Latent Dirichlet Allocation (LDA) [18] and Latent Semantic Index (LSI) [16]. Nonetheless, simpler methods have shown comparable performance while being more cost-efficient [23].
@@
In our work, we employ a BM25-based approach (e.g., Elasticsearch [42]) to collect the candidate source documents. Then, our subsequent modules combine the scalability of textual relevance with Transformer-based code understanding grounded in program semantics to retrieve buggy documents at top-ranked positions.
@@ 6.2 Query Reformulation
Studies by Mills et al. [85] and Rahman et al. [10] suggest that poor queries from bug reports may adversely affect the performance of a VSM-based search. Consequently, numerous approaches propose to construct queries to assist developers in their tasks [86, 11, 87, 88]. These studies for query construction fall into two categories: frequency-based and graph-based keyword selection methods [10].
In frequency-based methods, researchers have utilized TF-IDF [68] and its variants to extract meaningful keywords from both bug reports and source code for use as queries. For instance, Gay et al. [89] employed relevance feedback (RF) from users to update the query. They implemented Rocchio’s expansion [90] method, enhancing query performance by adding terms from relevant documents with increased weights and suppressing or removing terms from irrelevant documents. Haiduc et al. [87] proposed a technique that employs a machine learning model trained on 28 query properties to recommend the best reformulation strategy for a given query. These strategies include query reduction, Rocchio expansion, RSV expansion, or Dice expansion, selected based on the query’s properties and performance.
Graph-based methods analyze semantic and syntactic dependencies among words to determine their importance. Rahman and Roy [11] applied the PageRank algorithm [91] on constructed graphs to suggest search keywords from various sources. Subsequently, they explored genetic algorithms [92] for near-optimal search query construction, leveraging the vocabulary in bug reports for effective query building [10]. However, their genetic algorithm-based approach is costly and relies solely on the textual relevance between bug reports and source documents.
@@ 6.3 Deep Learning for Bug Localization
Recent advancements of deep learning in various domains (e.g., natural language processing, and machine translation) have encouraged its application to bug localization. DNNLOC, an early and influential work in the field of localization, is designed to identify potentially buggy documents by learning from multiple features (i.e., rVSM score, class name similarity, collaborative filtering, bug report recency, and bug report frequency) of bug reports and source code. However, its reliance on certain features like bug fixing recency and frequency is available for only 20-40% of the bugs [18]. So, the unavailability of these features might affect the performance of the model. A recent technique, FBL-BERT [93], incorporates the ColBERT model [94] for changeset-based bug localization. ColBERT uses a late interaction architecture, independently encoding bug reports and changesets with BERT, followed by efficient vector similarity calculations for relevance estimation. FBL-BERT enhances this by considering different levels of changeset granularity, enabling offline pre-computation of changeset embeddings, and employing a two-stage process of retrieving and reranking for bug localization. Nonetheless, the model may not be as effective in large-scale environments, where big software projects produces thousands of commits a day. Another recent bug localization technique, RLocator [31], is a first-of-its-kind reinforcement learning (RL) technique for bug localization. The technique attempts to optimize the ranking of a set of documents by employing RL agents based on an actor-critic [73] framework with entropy regularization, using a reward signal derived from ranking metrics (e.g., MAP and MRR) for localizing bugs. Other approaches, such as DeepLoator [95] proposed bug localization using Convolutional Neural Networks (CNN). DreamLoc [96] addresses the problem of bug localization by proposing a deep and wide architecture. The deep component applies attention mechanisms to capture textual relevance, while the wide component incorporates domain knowledge through a linear model using features such as bug-fixing history and code complexity. However, in addition to technical challenges, these techniques may encounter inefficiencies when dealing with large numbers of documents for bug localization.
@@
In contrast, our hybrid approach addresses these challenges by focusing on a limited set of documents using Transformer models for scalability and performance. Leveraging the self-attention mechanism of the Transformer model, we process both source code and bug reports simultaneously and analyze their semantics, which helps detect their contextual relevance. This methodology offers a promising avenue for overcoming the limitations faced by existing bug localization techniques.
@@ 7 Threats to Validity
Threats to internal validity relate to experimental errors and biases. Re-implementation of the existing baseline techniques could pose a threat. For BLUiR, Blizzard and RLocator, we collected the replication packages from Bench4BL, Rahman et al.’s GitHub repositories and from Zenodo [97, 19, 74]. However, since the Indri [67] library has become obsolete, we replaced that with Apache Lucene [63] in the BLUiR replication. For DNNLOC, as the replication package is not available by the original authors, we had to replicate it ourselves. While we acknowledge the potential for implementation errors, we addressed this concern by adhering to the settings and parameters of the original authors through extensive testing. We also repeat our experiments on two different datasets [11, 18] and compare the performance with baselines to mitigate any bias due to random trials.
Threats to external validity relate to the generalizability of our work. Even though IQLoc is evaluated using only Java code, the underlying models can be adapted to different programming languages through appropriate fine-tuning [98].
Threats to construct validity relate to the evaluation metrics for our work. We use several metrics such as Mean Average Precision (MAP), Mean Reciprocal Rank (MRR), and HIT@K, which are widely adopted in recommendation systems [11, 19, 99] and Information Retrieval [100, 101]. This confirms no or little threats to construct validity.
Finally, we adapted Bench4BL [23] dataset for our experiments, which might contain biases [102] and data quality issues (e.g., misclassified bugs, erroneously labeled buggy files). During the training of the Cross-Encoder model, we accepted method bodies containing buggy lines from Java classes as context. However, the impact of different context sizes was not well tested, which we consider as a scope of future work.
@@ 8 Conclusion and Future Work
Software bugs account for 50% of development time and cost billions of dollars annually. Despite significant research over the last few decades, bug localization remains a challenging task for developers. In this study, we introduce IQLoc, a novel hybrid approach that capitalizes on the strengths of both Information Retrieval (IR) and LLM-based program semantics understanding to support bug localization. Our approach enhances IR-based bug localization with reformulated queries, derived from the program understanding of Transformer models. By going beyond surface-level semantic relevance, IQLoc identifies buggy code more accurately. In our evaluation, we refined and extended the benchmark dataset Bench4BL and compared IQLoc against several baselines using three key metrics: Mean Average Precision (MAP), Mean Reciprocal Rank (MRR), and HIT@K. The results show that IQLoc consistently outperforms the baselines, with improvements of up to 60.59% in MAP, 64.58% in MRR, and 100.90% in HIT@K across two types of dataset splits. Additionally, further evaluation reveals that IQLoc achieves MAP improvements of 91.67% for bug reports with stack traces, 72.73% for those containing code elements, and 65.38% for those written solely in natural language. All these findings highlight the potential of IQLoc as a robust technique for bug localization that seamlessly integrates the textual relevance of traditional techniques with the program semantic understanding of modern Transformer-based models, setting a new benchmark for performance and reliability in addressing software bugs.
In the future, we plan to investigate how different representations of code structure (e.g., data flow, control flow [103]) and various tokenization strategies influence code understanding in bug localization using large language models (LLMs). Moreover, we aim to explore how agentic systems [104] can help address the multifaceted nature of bug reports in the context of bug localization using Information Retrieval (IR) techniques.

### v2

@@ 6.1 IR based Bug Localization
Bug localization has been a key area of research for decades, driven by the significant impacts and challenges of software bugs [practitioners_bug_localization_study]. There are two primary categories of bug localization: spectra-based and Information Retrieval (IR)-based [ir_bug_localization1]. Spectra-based methods are known for their complexity and limited scalability [spectra1, spectra2]. In our work, we primarily focus on Information Retrieval for bug localization.
@@
Traditional IR-based bug localization methods [saha_bleuir, ir_localization_lda_buglocator, ir_ml_amalgam, wang_amalgam+, blizzard] typically rely on the vector space model (VSM)[vector_space_model], which analyzes the token overlap between bug reports and source code to identify buggy documents. Several studies have extended VSM by integrating additional contexts, such as bug report history [saha2014effectiveness_bug_rep_history], code change history [wen2016locus], or version history [sisman2012incorporating], into the IR process. For instance, Zhou et al. introduced BugLocator [ir_localization_lda_buglocator] that employs a combined score of a modified VSM score and previous bug fix history to localize bugs. The traditional VSM-based scoring [vector_space_model] method tends to show bias towards longer documents. To address this, they modified the VSM-based scoring and introduced rVSM scoring, which enhances the computation of textual relevance between bug reports and code elements. BLUiR [saha_bleuir] captures four types of structural components from the source code (i.e., class names, method names, variable names, and comments) and two components from bug reports (i.e., bug title and bug description). It then forms pairwise combinations of these components to perform eight separate searches using Indri [Indri], leveraging a modified TF-IDF approach. The final localization score is computed by summing the scores obtained from these individual searches. AmaLgam [ir_ml_amalgam] integrates BLUiR and BugLocator techniques, along with version history inspired by Google by analyzing historical data from version control systems. These components undergo three independent systems to generate rankings before producing the final ranked lists. AmaLgam+ [wang_amalgam+] takes a step further by incorporating stack trace and bug reporter history, alongside AmaLgam’s contexts. It ranks files using five components and then returns a final ranked list of source documents for bug localization. Recently, an IR-based technique, PathIdea [chen2021pathidea], utilizes bug report logs (e.g., log snippets, stack traces) for bug localization. The authors employ a static analysis tool to construct a file-level call graph and reconstruct system execution paths from the logs. To determine the suspiciousness score for each file, they combine the VSM score, a log score that emphasizes files mentioned in the logs, and a path score that highlights files in the execution path. Some IR-based bug localization methods employ more complex mechanisms, such as Latent Dirichlet Allocation (LDA) [ir_localization_lda_buglocator] and Latent Semantic Index (LSI) [ir_bug_localization4]. Nonetheless, simpler methods have shown comparable performance while being more cost-efficient [bench4bl].
@@
In our work, we employ a BM25-based approach (e.g., Elasticsearch [elasticsearch]) to collect the candidate source documents. Then, our subsequent modules combine the scalability of textual relevance with Transformer-based code understanding grounded in program semantics to retrieve buggy documents at top-ranked positions.
@@ 6.2 Query Reformulation
Studies by Mills et al. [mills2020relationship] and Rahman et al. [rahman2021forgotten] suggest that poor queries from bug reports may adversely affect the performance of a VSM-based search. Consequently, numerous approaches propose to construct queries to assist developers in their tasks [query_reformulation_tosem, blizzard, haiduc2013automatic, bl_textrank_2]. These studies for query construction fall into two categories: frequency-based and graph-based keyword selection methods [rahman2021forgotten].
In frequency-based methods, researchers have utilized TF-IDF [tf-idf] and its variants to extract meaningful keywords from both bug reports and source code for use as queries. For instance, Gay et al. [gay2009use] employed relevance feedback (RF) from users to update the query. They implemented Rocchio’s expansion [rocchio_cambridge] method, enhancing query performance by adding terms from relevant documents with increased weights and suppressing or removing terms from irrelevant documents. Haiduc et al. [haiduc2013automatic] proposed a technique that employs a machine learning model trained on 28 query properties to recommend the best reformulation strategy for a given query. These strategies include query reduction, Rocchio expansion, RSV expansion, or Dice expansion, selected based on the query’s properties and performance.
Graph-based methods analyze semantic and syntactic dependencies among words to determine their importance. Rahman and Roy [blizzard] applied the PageRank algorithm [pagerank] on constructed graphs to suggest search keywords from various sources. Subsequently, they explored genetic algorithms [lambora2019genetic] for near-optimal search query construction, leveraging the vocabulary in bug reports for effective query building [rahman2021forgotten]. However, their genetic algorithm-based approach is costly and relies solely on the textual relevance between bug reports and source documents.
@@ 6.3 Deep Learning for Bug Localization
Recent advancements of deep learning in various domains (e.g., natural language processing, and machine translation) have encouraged its application to bug localization. DNNLOC, an early and influential work in the field of localization, is designed to identify potentially buggy documents by learning from multiple features (i.e., rVSM score, class name similarity, collaborative filtering, bug report recency, and bug report frequency) of bug reports and source code. However, its reliance on certain features like bug fixing recency and frequency is available for only 20-40% of the bugs [ir_localization_lda_buglocator]. So, the unavailability of these features might affect the performance of the model. A recent technique, FBL-BERT [ciborowska_fbl_bert], incorporates the ColBERT model [khattab2020colbert] for changeset-based bug localization. ColBERT uses a late interaction architecture, independently encoding bug reports and changesets with BERT, followed by efficient vector similarity calculations for relevance estimation. FBL-BERT enhances this by considering different levels of changeset granularity, enabling offline pre-computation of changeset embeddings, and employing a two-stage process of retrieving and reranking for bug localization. Nonetheless, the model may not be as effective in large-scale environments, where big software projects produce thousands of commits a day. Another recent bug localization technique, RLocator [RLocator], is a first-of-its-kind reinforcement learning (RL) technique for bug localization. The technique attempts to optimize the ranking of a set of documents by employing RL agents based on an actor-critic [konda2000actorcritic] framework with entropy regularization, using a reward signal derived from ranking metrics (e.g., MAP and MRR) for localizing bugs. Other approaches, such as DeepLocator [deeplocator] proposed bug localization using Convolutional Neural Networks (CNN). DreamLoc [dreamloc] addresses the problem of bug localization by proposing a deep and wide architecture. The deep component applies attention mechanisms to capture textual relevance, while the wide component incorporates domain knowledge through a linear model using features such as bug-fixing history and code complexity. However, in addition to technical challenges, these techniques may encounter inefficiencies when dealing with large numbers of documents for bug localization.
Recent work also investigates graph or tree-structured program representations for fault localization [AGFL, GNet4FL]. AGFL [AGFL] constructs Abstract Syntax Tree (AST) representation from source code and embeds its nodes using Word2Vec [word2vec]. It applies GraphSMOTE [graphSmote] for class balancing and then trains a GCN model [GCN_Graph_Conv_NN] with an attention mechanism to classify faulty nodes. GNet4FL [GNet4FL] preprocesses code by pruning AST nodes, embedding them with Word2Vec and t-SNE, and incorporating failed-test coverage as node features. It then applies GraphSAGE [GraphSAGE] for context-aware aggregation and uses an MLP to generate suspiciousness rankings for program statements. However, these techniques may underperform because generic Word2Vec embeddings fail to capture source-code semantics. Additionally, limited test coverage can reduce the availability of meaningful execution signals for many AST nodes (e.g., dynamically invoked code). DeepFL [DeepFL] integrates multiple dimensions of fault-diagnosis information (e.g., spectrum-based suspiciousness, mutation-based suspiciousness, code-complexity metrics, and textual similarity features) and learns how these feature patterns correspond to faulty code elements. By grouping related features hierarchically (e.g., merging mutation-based signals before combining with spectrum-based features), the model learns which combinations of signals reliably indicate a bug and produces a ranked list of suspicious methods. A recent approach [combining_error_guessing] combines developers’ error-guessing knowledge with the logical dependencies between methods. These signals are encoded as coverage matrices and fed into a CNN to predict and rank buggy methods. Another line of work explores multimodal bug-localization models, which integrate different modalities (e.g., textual bug reports, program-execution spectra, source code) within a unified learning framework. NetML [netml] integrates textual similarity, execution spectra, and suspicious-word features within a model that learns bug and method-specific weights. It further applies network-Lasso regularization on bug-report and method similarity graphs to encourage similar bugs and program elements to share parameters, thereby improving its ability to rank faulty methods. DeMoB [demob] adopts a similar multimodal perspective, encoding bug reports and source documents as distinct modalities using an attention-based BiLSTM and a multi-grained dynamic CNN–BiLSTM encoder, respectively. It detects buggy documents by projecting both representations into a shared embedding space and ranking source files according to their learned relevance to the bug report, optimized with a pairwise ranking loss. However, if training data is limited or execution spectra are unavailable, these techniques may struggle to learn alignments between bug reports and faulty code.
@@
In contrast, our hybrid approach addresses these challenges by focusing on a limited set of documents using Transformer models for scalability and performance. Leveraging the self-attention mechanism of the Transformer model, we process both source code and bug reports simultaneously and analyze their semantics, which helps detect their contextual relevance to buggy code elements. This methodology offers a promising avenue for overcoming the limitations of existing bug-localization techniques by moving beyond surface-level similarity and capturing deeper program-semantic understanding [program_uderstanding].
@@ 7 Threats to Validity
Threats to internal validity relate to experimental errors and biases. Re-implementation of the existing baseline techniques could pose a threat. For BLUiR [saha_bleuir], Blizzard [blizzard], BRTracer [BRTracer], BLIA [bl_code_change_BLIA] and RLocator [RLocator], we collected the replication packages from Bench4BL [bench4bl], Rahman et al.’s GitHub repositories [BLIZZARD_github] and from Zenodo [rlocator_replication_zenodo]. However, since the Indri [Indri] library has become obsolete, we replaced that with Apache Lucene [lucene] in the BLUiR replication. For DNNLOC [dnnloc_ir], LRBL [LRBL_learning2rank], LLmiRQ [LLmiRQ], as the replication package is not available, we had to replicate them ourselves. While we acknowledge the potential for implementation errors, we addressed this concern by adhering to the settings and parameters of the original authors and through extensive testing. We also repeat our experiments on two different datasets [blizzard, ir_localization_lda_buglocator] and compare the performance with baselines to mitigate any bias due to random trials. To further mitigate randomness and potential implementation bias, we repeated each random-split experiment five times and report averaged results.
Threats to external validity relate to the generalizability of our work. Even though IQLoc is evaluated using only Java code, the underlying models can be adapted to different programming languages through appropriate fine-tuning [domain_adapt].
Threats to construct validity relate to the evaluation metrics for our work. We use several metrics such as Mean Average Precision (MAP), Mean Reciprocal Rank (MRR), and HIT@K, which are widely adopted in recommendation systems [blizzard, saha_bleuir, Rack] and Information Retrieval [evaluation_ir, evaluation_metrics]. This confirms no or little threats to construct validity.
Finally, we adapted Bench4BL [bench4bl] dataset for our experiments, which might contain biases [bias_buglocalization] and data quality issues (e.g., misclassified bugs, erroneously labeled buggy files). During the training of the Cross-Encoder model, we accepted method bodies containing buggy lines from Java classes as context. However, the impact of different context sizes was not well tested, which we consider as a scope of future work.
@@ 8 Conclusion and Future Work
Software bugs account for 50% of development time and cost billions of dollars annually. Despite significant research over the last few decades, bug localization remains a challenging task for developers. In this study, we introduce IQLoc, a novel hybrid approach that capitalizes on the strengths of both Information Retrieval (IR) and LLM-based program semantics understanding to support bug localization. Our approach enhances IR-based bug localization with reformulated queries, derived from the program understanding of Transformer models. By going beyond surface-level semantic relevance, IQLoc identifies buggy code more accurately. In our evaluation, we refined and extended the benchmark dataset Bench4BL and compared IQLoc against several baselines using three key metrics: Mean Average Precision (MAP), Mean Reciprocal Rank (MRR), and HIT@K. The results show that IQLoc consistently outperforms the baseline techniques, with improvements of up to 100.40% in MAP, 64.58% in MRR, and 100.90% in HIT@K across two types of dataset splits. Further evaluation reveals that IQLoc achieves MAP improvements of 118.70% for bug reports with stack traces, 111.87% for those containing code elements, and 127.45% for those written solely in natural language. All these findings highlight the potential of IQLoc as a robust technique for bug localization that seamlessly integrates the textual relevance of traditional techniques with the program semantic understanding of modern Transformer-based models, setting a new benchmark for performance and reliability in addressing software bugs.
In the future, we plan to investigate how different representations of code structure (e.g., data flow, control flow [data_control_pdg_application]) and various tokenization strategies influence code understanding in bug localization using large language models (LLMs). Moreover, we aim to explore how agentic systems [swe-agent] can help address the multifaceted nature of bug reports in the context of bug localization using Information Retrieval (IR) techniques.